# Vortical waves in a quantum fluid with vector, axial and helical charges. II. Dissipative effects

### v1

@@ 3.2 Small $\tau_{H}$ limit: Recovering the CVW
Equations (55) and (56) of the previous subsection show that it is problematic to take the $k,\tau_{H}\rightarrow 0$ limit from the large-temperature expansion. In order to access the small $\tau_{H}\kappa_{\Omega}$ regime, we expand the quantity $\text{w}=\tau_{H}\omega$ and $\delta\mu_{\ell}$ in powers of $\tau_{H}$:
$\text{w}=\tau_{H}\omega=\text{w}_{0}+\text{w}_{1}\tau_{H}+\dots,\qquad\delta\mu_{\ell}=\delta\mu_{\ell;0}+\delta\mu_{\ell;1}\tau_{H}+\dots.$ (64)
It is convenient to work with w as in the $\tau_{H}\rightarrow 0$ limit, we can expect $\omega$ to have a $\tau_{H}^{-1}$ leading-order contribution, see e.g. Eq. (52). Additionally, from Eq. (45) we observe that w is a function of the combination $\kappa_{\Omega}\tau_{H}$ and consequently the small $\tau_{H}$ expansion naturally encapsulates the results for small wavenumbers.
@@
$\left(\frac{2\text{w}_{0}}{T^{2}}\right)^{2}\left(\sigma_{A}^{\omega}-\frac{T^{2}}{3}\Delta H\right)\left(\frac{i}{3}\sigma^{\omega}_{A}+\frac{2\text{w}_{0}}{T^{2}}[(\sigma^{\omega}_{A})^{2}-(\sigma^{\omega}_{H})^{2}]\right)=0,$ (65)
@@
The above corresponds to a purely dissipative helical mode, with $\text{w}_{h;0}=-i+O(\alpha_{V}^{2})$ at large temperature and $\text{w}_{h;0}\simeq-\frac{i\pi^{2}}{24}e^{|\alpha_{V}|}$ in the degenerate limit.
@@
$\frac{i}{3}\left[\left(\frac{2\text{w}^{\pm}_{{\mathfrak{a}};1}}{T^{2}}\right)^{2}\sigma_{A}^{\omega}\left(\sigma^{\omega}_{A}-\frac{T^{2}}{3}\Delta H\right)-\frac{A^{2}}{H}\kappa_{\Omega}^{2}\right]=0\Rightarrow\text{w}^{\pm}_{{\mathfrak{a}};1}=\pm\frac{AT^{2}\kappa_{\Omega}}{2\sigma^{\omega}_{A}\sqrt{H}}\left(1-\frac{T^{2}\Delta H}{3\sigma^{\omega}_{A}}\right)^{-1/2}.$ (67)
Using Eqs. (32) and (33), it can be seen that $\text{w}^{\pm}_{{\mathfrak{a}};1}=\pm 6\alpha_{V}\kappa_{\Omega}/(7\pi^{2})$ in the large-$T$ limit and ±$\pm 2\pi\kappa_{\Omega}/(\alpha_{V}^{2}\sqrt{3})$ in the degenerate limit. We remark that the two propagating modes $\text{w}^{\pm}_{{\mathfrak{a}};1}$ are necessarily different from the ones found in the small chemical potential expansion since those were not propagating at small wavenumber $k$. Therefore, we can anticipate that at finite but small chemical potential, two modes are propagating for small $k$. As the wavenumber $k$ increases, these modes will stop propagating, and we will have three purely dissipating modes until we reach yet another critical value of $k$, where the modes found in the small chemical potential limit start propagating again.
@@
$\frac{4\text{w}_{h;0}\text{w}_{h;1}}{T^{4}}\left(\sigma^{\omega}_{A}-\frac{T^{2}}{3}\Delta H\right)\left(\frac{6\text{w}_{h;0}}{T^{2}}[(\sigma_{A}^{\omega})^{2}-(\sigma_{H}^{\omega})^{2}]+\frac{2i}{3}\sigma^{\omega}_{A}\right)=0.$ (68)
@@
$\delta\mu_{A}^{{\mathfrak{a}};\pm}$ $\simeq\dfrac{AT^{2}\kappa_{\Omega}}{2\text{w}_{{\mathfrak{a}};1}^{\pm}\sigma_{A}^{\omega}}\delta\mu_{V}^{{\mathfrak{a}};\pm}=\sqrt{H}\left(1-\frac{T^{2}\Delta H}{3\sigma^{\omega}_{A}}\right)^{-1/2}\delta\mu_{V}^{{\mathfrak{a}};\pm}\,,$
@@
$\delta\mu_{V}^{h}\simeq\frac{T^{2}}{2\text{w}_{h;0}H\sigma^{\omega}_{A}}\frac{B\sigma^{\omega}_{A}-A\sigma^{\omega}_{H}}{\sigma_{A}^{\omega}-\frac{T^{2}}{3}\Delta H}\kappa_{\Omega}\tau_{H}\delta\mu^{h}_{H}\,,\qquad\delta\mu_{A}^{h}\simeq-\dfrac{\sigma_{H}^{\omega}}{\sigma_{A}^{\omega}}\delta\mu_{H}^{h}.$ (70)
@@
The Chiral Vortical Wave (CVW) can be recovered as follows. The traditional derivation of the CVW did not include fluctuations of the helical chemical potential. The freezing of the helical degree of freedom is naturally implemented if one requires instantaneous dissipation of helicity fluctuations, i.e. $\tau_{H}\to 0\,$. The angular frequencies $\omega_{h}=\text{w}_{h}/\tau_{H}$ and $\omega^{\pm}_{\mathfrak{a}}=\text{w}^{\pm}_{\mathfrak{a}}/\tau_{H}$ of the three modes found at small $\tau_{H}$ can be written in the leading order as:
@@
$\omega_{h}=-\frac{iT^{2}\sigma_{A}^{\omega}}{6\tau_{H}[(\sigma^{\omega}_{A})^{2}-(\sigma^{\omega}_{H})^{2}]}\,,\hskip 28.45274pt\omega^{\pm}_{\mathfrak{a}}=\pm\frac{AT^{2}}{2\sigma^{\omega}_{A}\sqrt{H}}\left(1-\frac{T^{2}\Delta H}{3\sigma^{\omega}_{A}}\right)^{-1/2}\frac{k\Omega}{T}\,.$ (71)
@@
Indeed, as $\tau_{H}\to 0$, we find that the helical mode dissipates very quickly, removing any fluctuation of helical chemical potential. In contrast, the axial modes give a propagating wave mixing only axial and vector degrees of freedom, with $\delta\mu_{H}=O(\tau_{H})$ [see Eq. (69)], i.e. the Chiral Vortical Wave Jiang:2015cva . Note that the CVW is only present at finite (vector) chemical potential $\mu_{V}$. In agreement with this property, the angular frequency of our axial modes, given by the second relation in Eq. (71), does not vanish provided $\mu_{V}\neq 0$ as it follows from the definition (31) of the coefficient $A$ and the absence of the leading $\mu_{V}$ behaviour in other terms in this expression.
@@
We remark that at finite (non-zero) $\tau_{H}$, the Chiral Vortical Wave (CVW) is propagating for small wavenumbers, but it can stop propagating at some finite wavenumber. These aspects will be further clarified in Sec. 3.4 (see also Fig. 4).
Refer to caption Refer to caption
@@
Figure 2: Profiles of the perturbations in the vector, axial and helical chemical potentials at time $c_{h}t/\sigma=5$ for the Gaussian initial conditions in Eq. (61), set by the initial Gaussian width $\sigma$ and the velocity $c_{h}$ of the helical vortical wave in a neutral, non-dissipating plasma (51), for various values of the helical relaxation time $\tau_{H}$. The helicity-preserving limit limit $\tau_{H}\rightarrow\infty$ corresponds to the helical vortical wave solutions shown in Eq. (63), while the helicity-frozen limit $\tau_{H}\rightarrow 0$ corresponds to the axial vortical wave described in Eq. (74). The parameters of the background state are $\mu_{V}=650$ MeV and $T=300$ MeV, while the rotation parameter is set to $\Omega=6.6$ MeV.
@@
$\bar{\mu}_{A}(t,z)$ $=\frac{\delta\mu_{V}}{2}\sqrt{H\left(1-\frac{T^{2}\Delta H}{3\sigma^{\omega}_{A}}\right)}\left[e^{-(z-v_{\mathfrak{a}}t)^{2}/2\sigma^{2}}-e^{-(z+v_{\mathfrak{a}}t)^{2}/2\sigma^{2}}\right].$ (74)
The above solution is confirmed in Fig. 2, where we took $T=300$ MeV, $\mu_{V}=650$ MeV and $\mu_{A}=\mu_{H}=0$ for the background state.
3.3 Large chemical potential limit
@@
$a(\alpha_{V})\omega^{3}+\dfrac{i}{\tau_{H}}b(\alpha_{V})\omega^{2}-\kappa^{2}_{\Omega}c(\alpha_{V})\omega-\dfrac{i}{\tau_{H}}\kappa_{\Omega}^{2}d(\alpha_{V})=0\,,$ (75)
@@
$a$ $=\frac{8}{T^{6}}\left(\sigma^{\omega}_{A}-\frac{T^{2}}{3}\Delta H\right)\left[(\sigma^{\omega}_{A})^{2}-(\sigma^{\omega}_{H})^{2}\right],$ $b$ $=\frac{4}{3T^{4}}\sigma_{A}^{\omega}\left(\sigma_{A}^{\omega}-\frac{T^{2}}{3}\Delta H\right),$
@@
$\lim_{|\alpha_{V}|\to\infty}a(\alpha_{V})$ $\to\dfrac{8\alpha_{V}^{4}}{3\pi^{6}}e^{-|\alpha_{V}|}\,,$ $\lim_{|\alpha_{V}|\to\infty}b(\alpha_{V})$ $\to\dfrac{\alpha_{V}^{4}}{9\pi^{4}}\,,$ (77)
$\lim_{|\alpha_{V}|\to\infty}c(\alpha_{V})$ $\to\dfrac{32}{9\pi^{4}}e^{-|\alpha_{V}|}\,,$ $\lim_{|\alpha_{V}|\to\infty}d(\alpha_{V})$ $=\dfrac{4}{27\pi^{2}}\,.$ (78)
Consequently, as $|\alpha_{V}|\to\infty$, the system hosts one purely imaginary mode and two purely propagating modes with real-valued energy dispersion relations, respectively:
$\omega_{\rm im.}=-\dfrac{i\pi^{2}}{24\tau_{H}}e^{|\alpha_{V}|}\,,\qquad\omega_{\pm}=\pm\dfrac{2\pi\Omega}{\sqrt{3}\alpha_{V}^{2}T}k\,,$ (79)
where we used the definition of $\kappa_{\Omega}$ in Eq. (28). From the above results, we find that as we increase the vector chemical potential, the non-propagating mode $\omega_{\rm im.}$ dissipates exponentially fast, whereas the phase speed $\omega_{\pm}/k=\pm 2\pi\Omega/\alpha_{V}^{2}\sqrt{3}$ of the other two modes vanishes quadratically, as $\alpha_{V}^{-2}$. This result is consistent with Fig. 5.
@@
Figure 3: Propagation of an initial Gaussian perturbation in the vector chemical potential, shown in Eq. (61), for the degenerate limit discussed in Sec. 3.3. The profiles are represented at three time instances, corresponding to $\tilde{v}t/\sigma=0$, $2.5$ and $5$, where the phase velocity of the propagating solution in the degenerate limit $\tilde{v}$ is given in Eq. (83). The background state has $\mu_{V}=200$ MeV and $T=10$ MeV. We took $\Omega=6.6$ MeV and $\tau_{H}=2\sigma/\tilde{v}$. We also represented the analytical solution for the limit $\alpha_{V}\rightarrow\infty$, shown in Eq. (84).
The general solution for $\delta\mu_{\chi}$ and $\delta\mu_{\tilde{\chi}}$ for arbitrary vector chemical potential and helical relaxation time is given by
$\delta\mu_{\chi}$ $=\frac{\kappa_{\Omega}T^{2}\left(AT^{2}\tau_{H}^{-1}-6i(A+s_{V}B)\omega_{*}(\sigma_{A}^{\omega}-s_{V}\sigma_{H}^{\omega})\right)}{2\omega_{*}\left(\sigma_{A}^{\omega}T^{2}\tau_{H}^{-1}-6i\omega_{*}\left[(\sigma_{A}^{\omega})^{2}-(\sigma_{H}^{\omega})^{2}\right]\right)}\delta\mu_{V}\,,$
$\delta\mu_{\tilde{\chi}}$ $=\frac{\kappa_{\Omega}T^{2}\left(AT^{2}\tau_{H}^{-1}-6i(A-s_{V}B)\omega_{*}(\sigma_{A}^{\omega}+s_{V}\sigma_{H}^{\omega})\right)}{2\omega_{*}\left(\sigma_{A}^{\omega}T^{2}\tau_{H}^{-1}-6i\omega_{*}\left[(\sigma_{A}^{\omega})^{2}-(\sigma_{H}^{\omega})^{2}\right]\right)}\delta\mu_{V}\,.$ (80)
From the previous expressions, it appears that the degenerate limit $\mu_{V}\to\infty$ and the helicity-conservation limit $\tau_{H}^{-1}\to 0$ do not commute. Naively, we could take $|\mu_{V}|/T\to\infty$ for some finite $\tau_{H}^{-1}$. In such case, the amplitudes for the propagating modes $\omega_{\pm}$ read
$\delta\mu_{\chi}^{\pm}$ $\simeq\delta\mu_{\tilde{\chi}}^{\pm}\simeq\frac{\kappa_{\Omega}T^{2}}{2\omega^{\pm}}\dfrac{A}{\sigma^{\omega}_{A}}\delta\mu_{V}^{\pm}\simeq\pm\dfrac{|\alpha_{V}|}{\sqrt{3}\pi}\delta\mu_{V}^{\pm}\,,$ (81)
@@
$\delta\mu_{\chi}^{\textrm{im.}}\simeq-\dfrac{2}{\alpha_{V}^{2}}e^{-|\alpha_{V}|}\delta\mu_{\tilde{\chi}}^{\textrm{im.}}\,\qquad\delta\mu_{V}^{\textrm{im.}}\simeq-\dfrac{72i\kappa_{\Omega}\tau_{H}}{\pi^{2}\alpha_{V}^{2}}e^{-2|\alpha_{V}|}\delta\mu_{\tilde{\chi}}^{\textrm{im.}}$ (82)
In Fig. 3, we show the evolution of the initial Gaussian profile of the vector chemical potential of Eq. (61) in the nearly-degenerate case when $\mu_{V}=200$ MeV and $T=10$ MeV. The plots show the profiles of $\delta\bar{\mu}_{V}(t,z)$, $\delta\bar{\mu}_{A}(t,z)$ and $\delta\bar{\mu}_{H}(t,z)$ at initial time and at times satisfying $\tilde{v}t/\sigma=2.5$ and $5$, where
$\tilde{v}=\frac{|{\rm Re}(\omega_{\pm})|}{k}=\frac{2\pi\Omega}{\alpha_{V}^{2}T\sqrt{3}}\,,$ (83)

### v2

@@ 3.2 Small $\tau_{H}$ limit: Recovering the CVW
Equations (51) and (52) of the previous subsection show that it is problematic to take the $k,\tau_{H}\rightarrow 0$ limit from the small chemical potential expansion. In order to access the small $\tau_{H}\kappa_{\Omega}$ regime, we expand the quantity $\text{w}=\tau_{H}\omega$ and $\delta\mu_{\ell}$ in powers of $\tau_{H}$:
$\text{w}=\tau_{H}\omega=\text{w}_{0}+\text{w}_{1}\tau_{H}+\dots,$
$\delta\mu_{\ell}=\delta\mu_{\ell;0}+\delta\mu_{\ell;1}\tau_{H}+\dots.$ (60)
It is convenient to work with w as in the $\tau_{H}\rightarrow 0$ limit, we can expect $\omega$ to have a $\tau_{H}^{-1}$ leading-order contribution, see e.g. Eq. (48). Additionally, from Eq. (41), we observe that w is a function of the combination $\kappa_{\Omega}\tau_{H}=k\Omega\tau_{H}/T$ [see Eq. (15b)] and consequently, the small $\tau_{H}$ expansion naturally encapsulates the results for small wavenumbers.
@@
$\left(\frac{2\text{w}_{0}}{T^{2}}\right)^{2}\left(\sigma_{A}^{\omega}-\frac{T^{2}}{3}\Delta H\right)$
×$\times\left(\frac{i}{3}\sigma^{\omega}_{A}+\frac{2\text{w}_{0}}{T^{2}}[(\sigma^{\omega}_{A})^{2}-(\sigma^{\omega}_{H})^{2}]\right)=0,$ (61)
@@
The above corresponds to a purely dissipative helical mode, with $\text{w}_{h;0}=-i+O(\alpha_{V}^{2})$ at large temperature and $\text{w}_{h;0}\simeq-\frac{i\pi^{2}}{24}e^{\lvert\alpha_{V}\rvert}$ in the degenerate limit.
@@
$\frac{i}{3}\left[\left(\frac{2\text{w}^{\pm}_{{\mathfrak{a}};1}}{T^{2}}\right)^{2}\sigma_{A}^{\omega}\left(\sigma^{\omega}_{A}-\frac{T^{2}}{3}\Delta H\right)-\frac{A^{2}}{H}\kappa_{\Omega}^{2}\right]=0$
$\Rightarrow\text{w}^{\pm}_{{\mathfrak{a}};1}=\pm\frac{AT^{2}\kappa_{\Omega}}{2\sigma^{\omega}_{A}\sqrt{H}}\left(1-\frac{T^{2}\Delta H}{3\sigma^{\omega}_{A}}\right)^{-1/2}.$ (63)
Using Eqs. (29) and (30), it can be seen that $\text{w}^{\pm}_{{\mathfrak{a}};1}=\pm 6\alpha_{V}\kappa_{\Omega}/(7\pi^{2})$ in the large-$T$ limit and ±$\pm 2\pi\kappa_{\Omega}/(\alpha_{V}^{2}\sqrt{3})$ in the degenerate limit. We remark that the two propagating modes $\text{w}^{\pm}_{{\mathfrak{a}};1}$ are necessarily different from the ones found in the small chemical potential expansion since those were not propagating at small wavenumber $k$. Therefore, we can anticipate that at finite but small chemical potential, two modes are propagating for small $k$. As the wavenumber $k$ increases, these modes will stop propagating, and we will have three purely dissipating modes until we reach yet another critical value of $k$, where the modes found in the small chemical potential limit start propagating again – see Sec. 3.4 for more details.
@@
$\frac{4\text{w}_{h;0}\text{w}_{h;1}}{T^{4}}\left(\sigma^{\omega}_{A}-\frac{T^{2}}{3}\Delta H\right)$
×$\times\left(\frac{6\text{w}_{h;0}}{T^{2}}[(\sigma_{A}^{\omega})^{2}-(\sigma_{H}^{\omega})^{2}]+\frac{2i}{3}\sigma^{\omega}_{A}\right)=0.$ (64)
@@
$\delta\mu_{A}^{{\mathfrak{a}};\pm}$ $\simeq\dfrac{AT^{2}\kappa_{\Omega}}{2\text{w}_{{\mathfrak{a}};1}^{\pm}\sigma_{A}^{\omega}}\delta\mu_{V}^{{\mathfrak{a}};\pm}$
=±$=\pm\sqrt{H}\left(1-\frac{T^{2}\Delta H}{3\sigma^{\omega}_{A}}\right)^{1/2}\delta\mu_{V}^{{\mathfrak{a}};\pm}\,,$
@@
$\delta\mu_{V}^{h}\simeq\frac{T^{2}}{2\text{w}_{h;0}H\sigma^{\omega}_{A}}\frac{B\sigma^{\omega}_{A}-A\sigma^{\omega}_{H}}{\sigma_{A}^{\omega}-\frac{T^{2}}{3}\Delta H}\kappa_{\Omega}\tau_{H}\delta\mu^{h}_{H}\,,$
$\delta\mu_{A}^{h}\simeq-\dfrac{\sigma_{H}^{\omega}}{\sigma_{A}^{\omega}}\delta\mu_{H}^{h}.$ (66)
@@
The Chiral Vortical Wave (CVW) can be recovered as follows. The traditional derivation of the CVW Jiang:2015cva did not include fluctuations of the helical chemical potential. The freezing of the helical degree of freedom is naturally implemented if one requires instantaneous dissipation of helicity fluctuations, i.e. $\tau_{H}\to 0\,$. The angular frequencies $\omega_{h}=\text{w}_{h}/\tau_{H}$ and $\omega^{\pm}_{\mathfrak{a}}=\text{w}^{\pm}_{\mathfrak{a}}/\tau_{H}$ of the three modes found at small $\tau_{H}$ can be written in the leading order as: [^4] The angular frequencies $\omega_{h}$ and $\omega^{\pm}_{\mathfrak{a}}$ in Eq. (67) of this article should not be confused with their non-dissipative counterparts, $\omega^{\pm}_{h}$ and $\omega_{\mathfrak{a}}$, obtained in, respectively, Eqs. (77) and (90) of the companion paper Morales-Tejera:2024uzg .
$\omega_{h}=-\frac{iT^{2}\sigma_{A}^{\omega}}{6\tau_{H}[(\sigma^{\omega}_{A})^{2}-(\sigma^{\omega}_{H})^{2}]}\,,$
@@
$\omega^{\pm}_{\mathfrak{a}}=\pm\frac{AT^{2}}{2\sigma^{\omega}_{A}\sqrt{H}}\left(1-\frac{T^{2}\Delta H}{3\sigma^{\omega}_{A}}\right)^{-1/2}\frac{k\Omega}{T}\,.$ (67)
@@
As $\tau_{H}\to 0$, we find that the helical mode dissipates very quickly, removing any fluctuation of helical chemical potential. In contrast, the axial modes give a propagating wave mixing only axial and vector degrees of freedom, with $\delta\mu_{H}=O(\tau_{H})$ [see Eq. (65)], i.e. the Chiral Vortical Wave Jiang:2015cva ; Gorbar:2017toh . In fact, the pair of modes in Eq. (67) coincides exactly with the result obtained in Ref. Gorbar:2017toh for the Chiral Vortical Wave. Using the definitions for $A$, $H$, $\Delta H$ and $\sigma_{A}$ given below Eq. (27), the pair of modes in Eq. (67) become
$\omega_{a}^{\pm}=\dfrac{6\pi T\mu_{V}(\pi^{2}T^{2}+5\mu_{V}^{2})k\Omega}{\sqrt{7\pi^{4}T^{4}+6\pi^{2}T^{2}\mu_{V}^{2}+15\mu_{V}^{4}}}$
×$\times\dfrac{1}{\sqrt{(\pi^{2}T^{2}+3\mu_{V}^{2})(7\pi^{4}T^{4}+30\pi^{2}T^{2}\mu_{V}^{2}+15\mu_{V}^{4})}}\,,$ (68)
which is the same as the modes obtained in Eq. (63) of Ref. Gorbar:2017toh upon substitution of Eq. (26) in the same paper. It is important to note that there is a discrepancy between Eq. (68) and the result of Ref. Jiang:2015cva , where the fluctuations in the energy-momentum sector were not taken into account (see Ref. Gorbar:2017toh for a detailed discussion).
Note that the CVW is only present at finite (vector) chemical potential $\mu_{V}$. In agreement with this property, the angular frequency of our axial modes, given by the second relation in Eq. (67), does not vanish provided $\mu_{V}\neq 0$ as it follows from the definition (28) of the coefficient $A$ and the absence of the leading $\mu_{V}$ behaviour in other terms in this expression.
@@
We remark that at finite (non-zero) $\tau_{H}$, the Chiral Vortical Wave (CVW) is propagating for small wavenumbers, but it can stop propagating when the wavenumber exceeds certain finite value. These aspects will be further clarified in Sec. 3.4 (see also Fig. 4).
@@
Figure 2: Profiles of the perturbations in the vector, axial and helical chemical potentials at time $c_{h}t/\sigma=5$ for the Gaussian initial conditions in Eq. (57), set by the initial Gaussian width $\sigma$ and the velocity $c_{h}$ of the helical vortical wave in a neutral plasma with conserved axial charge (47), for various values of the helical relaxation time $\tau_{H}$. The helicity-preserving limit $\tau_{H}\rightarrow\infty$ corresponds to the helical vortical wave solutions shown in Eq. (59), while the helicity-frozen limit $\tau_{H}\rightarrow 0$ corresponds to the axial vortical wave described in Eq. (71).We took $\alpha_{V}=13/6$ for the background state.
@@
$\bar{\mu}_{A}(t,z)$ $=\frac{\delta\mu_{V}}{2}\sqrt{H\left(1-\frac{T^{2}\Delta H}{3\sigma^{\omega}_{A}}\right)}$
×$\times\left[e^{-(z-v_{\mathfrak{a}}t)^{2}/2\sigma^{2}}-e^{-(z+v_{\mathfrak{a}}t)^{2}/2\sigma^{2}}\right].$ (71)
The above solution agrees with the $\tau_{H}\rightarrow 0$ limit shown in Fig. 2, where we took an unpolarized background state ($\mu_{A}=\mu_{H}=0$) with $\alpha_{V}=13/6\simeq 2.16$, corresponding to a ratio $c_{h}(\alpha_{V})/c_{a}(\alpha_{V})\simeq 2$ between the velocities of the pure helical-vortical and chiral-vortical waves. This figure also demonstrates the smooth transition from the HVW in the helicity-conserving limit to the CVW as $\tau_{H}$ is decreased from $\infty$ to 00.
3.3 Large vector chemical potential limit
@@
$a(\alpha_{V})\omega^{3}+\dfrac{i}{\tau_{H}}b(\alpha_{V})\omega^{2}-\kappa^{2}_{\Omega}c(\alpha_{V})\omega-\dfrac{i}{\tau_{H}}\kappa_{\Omega}^{2}d(\alpha_{V})\\ =0\,,$ (72)
@@
$a$ $=\frac{8}{T^{6}}\left(\sigma^{\omega}_{A}-\frac{T^{2}}{3}\Delta H\right)\left[(\sigma^{\omega}_{A})^{2}-(\sigma^{\omega}_{H})^{2}\right],$
$b$ $=\frac{4}{3T^{4}}\sigma_{A}^{\omega}\left(\sigma_{A}^{\omega}-\frac{T^{2}}{3}\Delta H\right),$
@@
$\lim_{\lvert\alpha_{V}\rvert\to\infty}a(\alpha_{V})$ $\to\dfrac{8\alpha_{V}^{4}}{3\pi^{6}}e^{-\lvert\alpha_{V}\rvert}\,,\lim_{\lvert\alpha_{V}\rvert\to\infty}b(\alpha_{V})\to\dfrac{\alpha_{V}^{4}}{9\pi^{4}}\,,$ (74)
$\lim_{\lvert\alpha_{V}\rvert\to\infty}c(\alpha_{V})$ $\to\dfrac{32}{9\pi^{4}}e^{-\lvert\alpha_{V}\rvert}\,,\lim_{\lvert\alpha_{V}\rvert\to\infty}d(\alpha_{V})=\dfrac{4}{27\pi^{2}}\,.$ (75)
Consequently, as $\lvert\alpha_{V}\rvert\to\infty$, the system hosts one purely imaginary mode and two purely propagating modes with real-valued energy dispersion relations, respectively:
$\omega_{\rm im.}=-\dfrac{i\pi^{2}}{24\tau_{H}}e^{\lvert\alpha_{V}\rvert}\,,\qquad\omega_{\pm}=\pm\dfrac{2\pi\Omega}{\alpha_{V}^{2}T\sqrt{3}}k\,,$ (76)
where we used the definition of $\kappa_{\Omega}$ in Eq. (15b). From the above results, we find that as we increase the vector chemical potential, the non-propagating mode $\omega_{\rm im.}$ dissipates exponentially fast, whereas the phase speed $\omega_{\pm}/k=\pm 2\pi\Omega/\alpha_{V}^{2}T\sqrt{3}$ of the other two modes vanishes quadratically, as $\alpha_{V}^{-2}$. This result is consistent with Fig. 5, which will be discussed in Sec. 3.4.
@@
Refer to caption Refer to caption
Figure 3: Propagation of an initial Gaussian perturbation in the vector chemical potential, shown in Eq. (57), for the degenerate limit discussed in Sec. 3.3. The profiles are represented at three time instances, corresponding to $\tilde{v}t/\sigma=0$, $2.5$ and $5$, where the phase velocity of the propagating solution in the degenerate limit $\tilde{v}$ is given in Eq. (80). The background state has $\alpha_{V}=20$ and $\tau_{H}=2\sigma/\tilde{v}$. We also represented the analytical solution for the limit $\alpha_{V}\rightarrow\infty$, shown in Eq. (81). The bottom right panel shows the dependence of $\delta\mu_{\tilde{\chi}}$ on $\tau_{H}$.
The general solution for $\delta\mu_{\chi}=\delta\mu_{A}+s_{V}\delta\mu_{H}$ and $\delta\mu_{\tilde{\chi}}=\delta\mu_{A}-s_{V}\delta\mu_{H}$ for arbitrary vector chemical potential $\mu_{V}$, where $s_{V}={\rm sgn}(\mu_{V})$, and arbitrary helical relaxation time, is given by
$\delta\mu_{\chi}=$
$\frac{\kappa_{\Omega}T^{2}\left(AT^{2}\tau_{H}^{-1}-6i(A+s_{V}B)\omega_{*}(\sigma_{A}^{\omega}-s_{V}\sigma_{H}^{\omega})\right)}{2\omega_{*}\left(\sigma_{A}^{\omega}T^{2}\tau_{H}^{-1}-6i\omega_{*}\left[(\sigma_{A}^{\omega})^{2}-(\sigma_{H}^{\omega})^{2}\right]\right)}\delta\mu_{V}\,,$
$\delta\mu_{\tilde{\chi}}=$
$\frac{\kappa_{\Omega}T^{2}\left(AT^{2}\tau_{H}^{-1}-6i(A-s_{V}B)\omega_{*}(\sigma_{A}^{\omega}+s_{V}\sigma_{H}^{\omega})\right)}{2\omega_{*}\left(\sigma_{A}^{\omega}T^{2}\tau_{H}^{-1}-6i\omega_{*}\left[(\sigma_{A}^{\omega})^{2}-(\sigma_{H}^{\omega})^{2}\right]\right)}\delta\mu_{V}\,.$ (77)
From the previous expressions, it appears that the degenerate limit $\mu_{V}\to\infty$ and the helicity-conservation limit $\tau_{H}^{-1}\to 0$ do not commute. Naively, we could take $\lvert\mu_{V}\rvert/T\to\infty$ for some finite $\tau_{H}^{-1}$. In such case, the amplitudes for the propagating modes $\omega_{\pm}$ read
$\delta\mu_{\chi}^{\pm}$ $\simeq\delta\mu_{\tilde{\chi}}^{\pm}\simeq\frac{\kappa_{\Omega}T^{2}}{2\omega^{\pm}}\dfrac{A}{\sigma^{\omega}_{A}}\delta\mu_{V}^{\pm}\simeq\pm\dfrac{\lvert\alpha_{V}\rvert}{\pi\sqrt{3}}\delta\mu_{V}^{\pm}\,,$ (78)
@@
$\delta\mu_{\chi}^{\textrm{im.}}$ $\simeq-\dfrac{2}{\alpha_{V}^{2}}e^{-\lvert\alpha_{V}\rvert}\delta\mu_{\tilde{\chi}}^{\textrm{im.}}\,,$
$\delta\mu_{V}^{\textrm{im.}}$ $\simeq-\dfrac{72i\kappa_{\Omega}\tau_{H}}{\pi^{2}\alpha_{V}^{2}}e^{-2\lvert\alpha_{V}\rvert}\delta\mu_{\tilde{\chi}}^{\textrm{im.}}.$ (79)
In Fig. 3, we show the evolution of the initial Gaussian profile of the vector chemical potential of Eq. (57) in the nearly-degenerate case when $\alpha_{V}=20$. The plots show the profiles of $\delta\bar{\mu}_{V}(t,z)$, $\delta\bar{\mu}_{A}(t,z)$ and $\delta\bar{\mu}_{H}(t,z)$ at initial time and at times satisfying $\tilde{v}t/\sigma=2.5$ and $5$, where
$\tilde{v}=\frac{\lvert{\rm Re}(\omega_{\pm})\rvert}{k}=\frac{2\pi\Omega}{\alpha_{V}^{2}T\sqrt{3}}\,,$ (80)